# The Double-Burst Nature and Early Afterglow Evolution of Long GRB 110801A

### v1

@@ I Introduction
Gamma-ray bursts (GRBs) are the most luminous explosions that originated from either the collapse of massive stars (Piran, 2005; Woosley and Bloom, 2006) or the merger of compact binaries (Abbott et al., 2017). Following a catastrophic event, an outflow is powered by the central engine and is likely collimated into a relativistic jet during propagation (Salafia and Ghirlanda, 2022). The nature of prompt emission can be explained by the energy dissipation in the internal shock (Rees and Meszaros, 1994; Kobayashi et al., 1997; Elliott et al., 2014) or magnetic dissipation (Beniamini and Granot, 2016; Granot, 2016; Troja et al., 2017), although its complex radiation mechanisms remain poorly understood. The most natural candidate is synchrotron radiation generated from energetic electrons. However, the theoretical synchrotron spectrum predicts photon indices in lower energy band expected for the fast cooling synchrotron model are softer than typical photon indices (), which is the so-called fast cooling problem(Ghisellini et al., 2000). To reconcile this inconsistency, several modifications to the canonical synchrotron framework have been suggested, such as additional inverse Compton scattering (Nakar et al., 2009; Daigne et al., 2011) or a decaying magnetic field (Pe’er and Zhang, 2006; Uhm and Zhang, 2014; Zhao et al., 2014). Another alternative interpretation is the presence of a sub-dominant quasi-thermal component (in addition to a non-thermal component)(Guiriec et al., 2011; Peng et al., 2014; Valan and Larsson, 2021; Wang et al., 2025b). The afterglow is powered by synchrotron radiation from the external forward shock or reverse shock, produced as the jet interacts with the circumburst medium (Mészáros and Rees, 1997, 1999; Sari and Piran, 1999; Blake et al., 2005; Gendre et al., 2012; Arimoto et al., 2024; Wang et al., 2025a).
The prompt emission of GRB 110801A in -ray shows two comparable energetic episodes separated by a quiescence time of roughly 200 s like GRB 240529A (Sun et al., 2024). The optical observation starts during the -ray’s quiescent time and shows a sharp rise. The two isolated -ray episodes and the sharp transition in the optical rise imply the presence of ”refreshed shock”. The ”refreshed shock” scenario is commonly invoked to explain the re-brightening behavior observed in late-time afterglows (Lamb et al., 2020; Moss et al., 2023; Schroeder et al., 2024). In this paper, we report the analysis of long GRB 110801A data obtained by the Neil Gehrels Swift (Gehrels et al., 2004). Similar to Jin et al. (2023) and Zhou et al. (2023), we split the the ultraviolet/optical telescope (UVOT, Roming et al., 2005) data taken in the event mode to short time bins to explore the optical temporal behavior at the early stage of the burst. We conclude the optical flux is dominated by the earlier shock’s afterglow emission, while the X-ray and -ray flux is dominated by the later shell’s prompt emission.
@@ II.2 Data reduction
The data from Swift BAT, XRT and UVOT, were reduced with the HEASoft 6.35. The BAT light curve (15 - 350 keV) from s to s was extracted with a time bin size of 1 s and generated using standard procedures described in the Swift/BAT software guide111https://www.swift.ac.uk/analysis/bat/index.php. The XRT source and background spectral files in both PC and WT modes the redistribution matrix, and ancillary response files were reduced with the online Swift/XRT data products generator222https://www.swift.ac.uk/burst_analyser/(Evans et al., 2007, 2009). UVOT observed GRB 110801A in the V, B, U, W1 and WHITE bands for several epochs. For data taken in the image mode, we started from the level 2 UVOT products and made photometry with the standard aperture. For the WHITE-band and U-band data taken in the event mode, a fixed time bin of 50 s (20 s) were applied to extract the light curve in the WHITE (U) band. The standard aperture is also adopted to make photometries of the event data. The UVOT photometries are listed in Table LABEL:table:2
@@ IV.2 The spectral analysis of the afterglow
During the XSPEC fitting, the and models are set the same as mentioned above. We tried both simple power-law and cut-off power-law model; however, we cannot use the cut-off power-law model here because the cutoff is in the -ray band and we do not have that detection at this epoch.
@@ IV.3 Afterglow modeling
The interaction between the relativistic outflow of the GRB and the circumburst medium induces the formation of a pair of relativistic shocks, i.e., FS and RS components. Initially, the light curve is dominated by the RS emission characterized by a steep rise. The FS takes over soon after the U-band peak, and dominates the decay phase. The electrons accelerated by these shocks are capable of generating multi-band emissions via synchrotron radiation and inverse Compton scattering. Numerous computational paradigms for modeling afterglows have been developed, we use a Python-wrapped FORTRAN package444https://github.com/mikuru1096/ASGARD_GRBAfterglow(Ren et al., 2024) to perform numerical calculations of the radiation contributed from the FS component under the various effects. The calculation of the contribution from the RS in the early afterglow phase is referenced from (Yi et al., 2013). Therefore, in this model, the flux density at a certain time and frequency is:
@@
We utilize emcee (Foreman-Mackey et al., 2013) to conduct Bayesian inference on the parameters of the afterglow model (Eq. 7). The prior ranges for these parameters are provided in Table 3. We consider a -ray radiation efficiency (the minimum value reported in the statistical study by Zhang et al. (2007)), which corresponds to an upper bound of 55 for the prior of . Additionally, we define a log-likelihood term for the each data point, which corresponds to an observation at time in a band with central frequency , as follows:
where is a free parameter to characterize the systematic error. Furthermore, considering the difficulty in accurately calculating the extinction for the WHITE, U, and W1 bands, we introduced a free parameter to the magnitudes of each of these three bands. In the double-burst scenario, the first ejecta initially propagates through the medium while decelerating. The second ejecta, launched approximately during the active phase of the second pulse, moves at a constant velocity because the first ejecta has already swept up most of the ambient material. Consequently, it eventually catches up and collides with the first ejecta. By interpolating the numerical calculation results for the radius evolution of the first ejecta, we derive that the second ejecta, moving at a constant velocity, would catch up with the first ejecta approximately 370 s in the observer frame. We model the energy extraction from the catch-up result as a continuous energy injection in the form of , where is the energy injection index (Dai and Lu, 1998; Zhang and Mészáros, 2001; Zhang et al., 2006). Here we adopt the radiation efficiency of (), and as typical values. The best simultaneous fit results are shown in Figure 4 and the parameters are summarized in Figure 5. By fitting the light curves, we are able to constrain various parameters for both the FS and RS components. According to the FS and RS models, the modeled light curves reproduce both the multi-band optical data and the late-time X-ray emission well, with all derived parameters constrained within typical ranges. This consistency corroborates the hypothesis suggested by our joint spectral fitting that the optical emission is dominated by the afterglow contribution of the first burst. Furthermore, the initial Lorentz factor , the half-opening angle and the isotropic kinetic energy erg. The fit result constrains well the fraction of the forward shock energy converted into the energy of magnetic field and the reverse shock energy converted into the energy of magnetic field .

### v2

@@ I Introduction
Gamma-ray bursts (GRBs) represent the most energetic explosions in the universe that originated from either the collapse of massive stars (Piran, 2005; Woosley and Bloom, 2006) or the merger of compact binaries (Abbott et al., 2017). Following a catastrophic event, an outflow is powered by the central engine and is likely collimated into a relativistic jet during propagation (Salafia and Ghirlanda, 2022). The nature of prompt emission can be explained by the energy dissipation in the internal shock (Rees and Meszaros, 1994; Kobayashi et al., 1997; Elliott et al., 2014) or magnetic dissipation (Beniamini and Granot, 2016; Granot, 2016; Troja et al., 2017), although its complex radiation mechanisms remain poorly understood. The most natural candidate is synchrotron radiation generated from energetic electrons. However, the theoretical synchrotron spectrum predicts photon indices in lower energy band expected for the fast cooling synchrotron model are softer than typical photon indices (), which is the so-called fast cooling problem(Ghisellini et al., 2000). To reconcile this inconsistency, several modifications to the canonical synchrotron framework have been suggested, such as additional inverse Compton scattering (Nakar et al., 2009; Daigne et al., 2011) or a decaying magnetic field (Pe’er and Zhang, 2006; Uhm and Zhang, 2014; Zhao et al., 2014). Another alternative interpretation is the presence of a sub-dominant quasi-thermal component (in addition to a non-thermal component)(Guiriec et al., 2011; Peng et al., 2014; Valan and Larsson, 2021; Wang et al., 2025b). The afterglow is powered by synchrotron radiation from the external forward shock or reverse shock, produced as the jet interacts with the circumburst medium (Mészáros and Rees, 1997, 1999; Sari and Piran, 1999; Blake et al., 2005; Gendre et al., 2012; Arimoto et al., 2024; Wang et al., 2025a).
The prompt emission of GRB 110801A in -ray shows two comparable energetic episodes separated by a quiescence time of roughly 200 s like GRB 240529A (Sun et al., 2024). The optical observation starts during the -ray’s quiescent time and shows a sharp rise. The two isolated -ray episodes and the sharp transition in the optical rise imply the presence of ”refreshed shock”. The ”refreshed shock” scenario is commonly invoked to explain the re-brightening behavior observed in late-time afterglows (Lamb et al., 2020; Moss et al., 2023; Schroeder et al., 2024). In this paper, we report the analysis of long GRB 110801A data obtained by the Neil Gehrels Swift (Gehrels et al., 2004). Following the approach of Jin et al. (2023) and Zhou et al. (2023), we split the ultraviolet/optical telescope (UVOT, Roming et al., 2005) data taken in the event mode into finer time bins. This allowed us to investigate the optical temporal behavior during the early stage of the burst. We conclude the optical flux is dominated by the earlier shock’s afterglow emission, while the X-ray and -ray flux is dominated by the later shell’s prompt emission.
@@ II.2 Data reduction
We employed the HEASoft software package (version 6.35) to reduce the data obtained fromSwift BAT, XRT and UVOT. The BAT light curve (15 - 350 keV) from s to s was extracted with a time bin size of 1 s and generated using standard procedures described in the Swift/BAT software guide111https://www.swift.ac.uk/analysis/bat/index.php. The XRT source and background spectral files in both PC and WT modes the redistribution matrix, and ancillary response files were reduced with the online Swift/XRT data products generator222https://www.swift.ac.uk/burst_analyser/(Evans et al., 2007, 2009). UVOT observed GRB 110801A in the V, B, U, W1 and WHITE bands for several epochs. For data taken in the image mode, we started from the level 2 UVOT products and made photometry with the standard aperture. For the WHITE-band and U-band data taken in the event mode, a fixed time bin of 50 s (20 s) were applied to extract the light curve in the WHITE (U) band. The standard aperture is also adopted to make photometries of the event data. The UVOT photometries are listed in Table LABEL:table:2
@@ IV.2 The spectral analysis of the afterglow
During the XSPEC fitting, the and models are set the same as mentioned above. We attempted to fit the data with both SPL and CPL model; however, we cannot use the CPL model here because the cutoff is above the X-ray band and we do not have -ray band detection at this epoch.
@@ IV.3 Afterglow modeling
As the relativistic outflow of the GRB propagates into the circumburst medium, it drives a pair of relativistic shocks, i.e., FS and RS components. Initially, the light curve is dominated by the FS emission characterized by an onset in the WHITE band. The RS soon takes over, characterized by a steep rise in the U band. After the peak in the U band, FS dominates the decay phase. The electrons accelerated by relativistic shocks are capable of producing multi-band emissions via synchrotron radiation and inverse Compton scattering. Numerous computational paradigms for modeling afterglows have been developed. To perform numerical calculations of the radiation contributed from the FS component under the various effects, we use a Python-wrapped FORTRAN package444https://github.com/mikuru1096/ASGARD_GRBAfterglow(Ren et al., 2024). For the calculation of the contribution from the RS in the early afterglow phase, it is referenced from (Yi et al., 2013). Therefore, in this model, the flux density at a certain time and frequency is:
@@
We utilize emcee (Foreman-Mackey et al., 2013) to conduct Bayesian inference on the parameters of the afterglow model (Eq. 7). The prior ranges for these parameters are provided in Table 3. We consider a -ray radiation efficiency (the minimum value reported in the statistical study by Zhang et al. (2007)), which corresponds to an upper bound of 55 for the prior of . Furthermore, we introduce a log-likelihood term. For a given data point at time in a band with central frequency , the log-likelihood is calculated:
where is a free parameter to characterize the systematic error. Furthermore, considering the difficulty in accurately calculating the extinction for the WHITE, U, and W1 bands, we introduced a free parameter to the magnitudes of each of these three bands. In the double-burst scenario, the first ejecta initially propagate through the medium while decelerating. The second ejecta, launched approximately during the active phase of the second pulse, moves at a constant velocity because the first ejecta has already swept up most of the ambient material. Consequently, it eventually catches up and collides with the first ejecta. By interpolating the numerical calculation results for the radius evolution of the first ejecta, we derive that the second ejecta, moving at a constant velocity, would catch up with the first ejecta approximately 370 s in the observer frame. We model the energy extraction from the catch-up result as a continuous energy injection in the form of , where is the energy injection index (Dai and Lu, 1998; Zhang and Mészáros, 2001; Zhang et al., 2006). Here we adopt the radiation efficiency of (), and as typical values. The best simultaneous fit results are shown in Figure 4 and the parameters are summarized in Figure 5. By fitting the light curves, we are able to constrain various parameters for both the FS and RS components. According to the FS and RS models, the modeled light curves reproduce both the multi-band optical data and the late-time X-ray emission well, with all derived parameters constrained within typical ranges. This consistency corroborates the hypothesis suggested by our joint spectral fitting that the optical emission is dominated by the afterglow contribution of the first burst. Furthermore, the initial Lorentz factor , the half-opening angle and the isotropic kinetic energy erg. The fit result constrains well the fraction of the forward shock energy converted into the energy of magnetic field and the reverse shock energy converted into the energy of magnetic field .